Grade 11Mechanics

When the system is released from rest acceleration of block 2 and 3 w.r.t 1 is 1 m / s ^ 2 downward and 5m/s^2 downward acceleration of block 3 relative to 4 is 0 find absolute accelerations of blocks

padmapriya
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
a2 – a1 = 1
a3 – a1 = 5
a3 – a4 = 0 
Now realtive accl. b/w a3 and a4 is zero therfeore a3 = a4 = g
put in the eqns. and get the values.